2. A solution is prepared by dissolving 7.00 g of glycerin (C3H8O3) in 201 g of ethanol (C2H5OH). The freezing point of the solution is __________°C. The freezing point of pure ethanol is -114.6 °C at 1 atm. The molal-freezing-point-depression constant (Kf) for ethanol is 1.88 °C/m. The molar masses of glycerin and of ethanol are 92.1 g/mol and 46.1 g/mol, respectively.
Answer #1

dTf = Kf*m
molality = moles / kg of solvent
moles of glycerine = 7.00 / 92.1 = 0.0760 mol
molality = 0.0760 / 0.201 = 0.378 mol/kg
?Tf = 1.99*0.378 = 0.75